• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 20:51:46 on January23,2002

    Entry number 80428

    keyword=Temperature sensors cable fixed. In beam permit now

    Chris and MIke Seely fixed the cable problem for the CERNOX temperature sensors.
    We are all out and in beam permit now.